# Deep Backtracking Counterfactuals for Causally Compliant Explanations

### v3

@@ 1 Introduction
The present work further serves as a bridge between causal modeling and practical methods in the field of counterfactual explanations (Wachter et al., 2017; Beckers, 2022). As a causally grounded approach applicable to high-dimensional data, our method fills a gap in the existing literature between non-causal explanation tools, built for complex data such as images (e.g., Goyal et al., 2019; Boreiko et al., 2022), and causal methods that have only been applied to simple, low-dimensional settings (Bynum et al., 2024; von Kügelgen et al., 2023b).
@@ 2.1 Structural Causal Models
Let $\mathbf{X}=(X_{1},X_{2},...,X_{n})$ be a collection of potentially high-dimensional observable “endogenous” random variables. For instance, these variables could be high-dimensional objects such as images (e.g., the MNIST image in Fig. 1) or scalar feature variables (such as $t$ and $i$ in Fig. 1). The causal relationships among the $X_{i}$ are specified by a directed acyclic graph $G$ that is known. A structural causal model (Pearl, 2009) is characterized by a collection of structural equations $X_{i}\leftarrow f_{i}(\mathbf{X}_{\text{pa}(i)},U_{i})$, for $i=1,2,...,n,$ where $\mathbf{X}_{\text{pa}(i)}$ are the causal parents of $X_{i}$ as specified by $G$ and $\mathbf{U}=(U_{1},U_{2},...,U_{n})$ are exogenous latent variables. The acyclicity of $G$ ensures that for all $i$, we can recursively solve for $X_{i}$ to obtain a deterministic expression in terms of $\mathbf{U}$. Thus, there exists a unique function that maps $\mathbf{U}$ to $\mathbf{X}$, which we denote by $\mathbf{F}$,
@@ 2.2 Deep Invertible Structural Causal Models
In this work, we make the simplifying assumption that $f_{i}(\mathbf{x}_{\text{pa}(i)},\,\cdot\,)$ is invertible for any fixed $\mathbf{x}_{\text{pa}(i)}$111also known as bijective generation mechanism (see, e.g., Nasr-Esfahany et al., 2023). such that we can write
@@
$\mathbf{U}=\mathbf{F}^{-1}(\mathbf{X}).$
@@
(Rezende & Mohamed, 2015; Winkler et al., 2019) are constructed as a composition of invertible functions, hence rendering the entire function $f_{i}$ invertible in $u_{i}$. In addition, they are chosen such that the determinant of the Jacobian can be compted efficiently. These two attributes facilitate efficient training of $f_{i}$ via maximum likelihood.
@@ 2.3 Interventional and Backtracking Counterfactuals
leave all structural assignments unchanged. In order to set the antecedent $\mathbf{x}^{*}_{S}\neq\mathbf{x}_{S}$ true, they trace differences to the factual realization back to (ideally small) changes in the latent variables $\mathbf{U}$. These modified latent variables are represented by a new collection of variables $\mathbf{U}^{*}$ that depend on $\mathbf{U}$ via a backtracking conditional $p^{B}(\mathbf{u}^{\prime}\,|\,\mathbf{u})$ (red in Fig. 2(von Kügelgen et al., 2023b), which represents a probability density for computing similarity between $\mathbf{u}$ and $\mathbf{u}^{\prime}$ and which we assume to be decomposable, or factorized:
@@
$p^{B}_{i}(u^{\prime}_{i}\,|\,u_{i})\;\propto\;\text{exp}\left\{-d_{i}(u^{\prime}_{i},u_{i})\right\},$
where the $d_{i}$ are differentiable with respect to $u^{\prime}_{i}$. As we will see in the following, $d_{i}(u^{\prime}_{i},u_{i})$ can be interpreted as a distance function.
@@ 2.4 Counterfactual Explanations
where $d_{o}$ represents a distance function between observed variables. In the present work, we generally refer to methods implementing a variant of 2 as counterfactual explanations. We stress that $\hat{Y}$ is the prediction of a model ($f_{\hat{Y}}$) and thus always an effect of $X$. In general, the structural assignment $f_{\hat{Y}}$ is different from $f_{Y}$ (the assignment of the true variable $Y$ that is not predicted). For instance, $Y$ might be a cause of $X$ or might be confounded with $X$. We revisit the difference between $Y$ and $\hat{Y}$ in Section § 3.2.
@@ 3.1.1 Objective for Stochastic DeepBC
We sample from a distribution over counterfactuals $\mathbf{X}^{*}\,|\,\mathbf{x}^{*}_{S},\mathbf{x}$ for the factual realization $\mathbf{x}$, antecedent $\mathbf{x}^{*}_{S}$ and (known) backtracking conditional $p^{B}$. This distribution is characterized by the density
@@
where $\delta_{\mathbf{x}^{*}_{S}}(\,\cdot\,)$ refers to the dirac delta at $\mathbf{x}^{*}_{S}$. Intuitively, we can understand this distribution as describing counterfactuals that are likely given $\mathbf{x}$ in terms of latent components ($p^{B}$), while fullfilling the constraint of being compliant both with the antecedent $\mathbf{x}^{*}_{S}$ ($\delta_{\mathbf{x}^{*}_{S}}$) and with the causal laws $\mathbf{F}$. We further note that 3 is equivalent to a sampling problem within the structured latent space, i.e.,
@@
The distribution of $\mathbf{X}^{*}\,|\,\mathbf{x}^{*}_{S},\mathbf{x}$ in 3 is obtained as the push-forward of 4 by $\mathbf{F}$, for $\mathbf{x}=\mathbf{F}(\mathbf{u})$.
@@ 3.1.2 Objective for Mode DeepBC
We compute the mode of $p(\mathbf{x}^{\prime}\,|\,\mathbf{x}^{*}_{S},\mathbf{x})$ in 3, i.e., a single “most likely” counterfactual $\mathbf{x}^{*}$ for the factual realization $\mathbf{x}$ as a solution to the following constrained optimization problem:
@@
Intuitively, we can understand this optimization as finding a solution $\mathbf{x}^{*}$ that is close to the factual realization $\mathbf{x}$ in terms of its latent components. This situation is visualized on the Morpho-MNIST example in Fig. 1. We further note that 5 is equivalent to an optimization problem within the structured latent space,
@@
We obtain the solution of 5 by inserting the solution of 6 into $\mathbf{F}$.
@@ 3.2 Relation to Counterfactual Explanations
We can recover counterfactual explanations (§ 2.4) as a special form of mode DeepBC 5. To this end, we assume access to two variables with the following structural equations
@@
where we note that $\hat{Y}$ is not subject to additional randomness $U_{\hat{Y}}$. In this specific case, we observe that the mode DeepBC optimization problem 5 reduces to
@@
which can be interpreted as an instance of 2, where distance is measured in an unstructured latent space, governed by $f_{X}$. From this viewpoint, we can interpret DeepBC as a general form of counterfactual explanations 2 in two ways: Firstly, it accommodates non-deterministic relations among variables, taking into account the influence of noise on all variables. In the aforementioned instance 7, this can be modeled by $\hat{Y}\leftarrow f_{\hat{Y}}(X,U_{\hat{Y}})$ (demonstrated in Fig. 12 (b) of § D.2). Secondly, DeepBC accounts for multiple variables with more general underlying causal relationships. For example, there could be a third variable $Z$ related to both $X$ and $Y$ in 8 that could be modelled as well, as depicted in Fig. 3 (a). Rather than treating $Z$ as a new dimension in the predictor input $f_{\hat{Y}}$ in 7 (resulting in $f_{\hat{Y}}(X,Z)$, see Fig. 3 (b)), DeepBC allows for explicitly modeling the true causal relations between $Z$, $X$ and $Y$ (see Fig. 3 (a)) via (deep) structural equations, as outlined in Section § 2.1. The benefit of DeepBC is that it intends to answer queries regarding the true underlying variables (see query in Fig. 3 (a)) rather than being confined to model predictions that generally do not model the causal mechanisms according to the true relationships (see query in Fig. 3 (b)). We elaborate further on the benefits of our approach in Section § 3.5.
@@
In the next section (§ 3.3), we propose algorithms to generate backtracking counterfactuals in practice, based on the formulations presented in 3 and 5.
@@ 3.3 Algorithms
We rely on a penalty formulation to approximate 4 and 6 in order to account for the dirac measure and the constraint, respectively. Specifically, we consider the following energy (loss) function with respect to $\mathbf{u}^{\prime}$:
@@ 3.3.1 Algorithm for Stochastic DeepBC
and hence we can use 10 to generate approximate samples from the desired distribution 4. In practice, we apply an Euler-Maruyama discretization (Sauer, 2013) of 10. It is given as
@@ 3.3.2 Algorithm for Mode DeepBC
We approximately compute the mode of $\mathbf{X}^{*}\,|\,\mathbf{x}^{*}_{S},\mathbf{x}$ by directly minimizing the energy $\mathcal{L}(\mathbf{u}^{\prime};\,\mathbf{u},\mathbf{x}^{*}_{S})$ 9 with respect to $\mathbf{u}^{\prime}$, where $\mathbf{u}=\mathbf{F}^{-1}(\mathbf{x})$. Rather than performing gradient descent on the original objective, we empirically observe that using a first-order Taylor approximation of $\mathbf{F}_{S}$ at $\bar{\mathbf{u}}$ is beneficial when minimizing the distance
@@
where $\bar{\mathbf{u}}$ is the linearization point and $\mathbf{J}_{S}(\bar{\mathbf{u}})\coloneqq\nabla_{\mathbf{u}}\mathbf{F}_{S}(\bar{\mathbf{u}})^{\top}$ denotes the Jacobian matrix. As a result of this approximation, 9 is a convex quadratic function in $\mathbf{u}^{\prime}$ and can therefore be solved for its minimum $\hat{\mathbf{u}}^{*}$ in closed form:
@@
and $\mathbf{W}\coloneqq\text{diag}(w_{i})$ is a diagonal matrix containing the distance weights $w_{i}$. By default, we set $w_{i}=1$, for all $i$. A different choice of weights may be useful to encode the notion that certain noise variables are more stable (where large $w_{i}$ corresponds to a high stability) due to application-specific requirements. A detailed derivation of 12 is provided in § A.3.
Solving 12 once, starting from the initial condition $\bar{\mathbf{u}}=\mathbf{u}$, does not accurately fulfill the constraint due to the constraint linearization, except for special cases. We thus apply an iterative algorithm similar to the Levenberg-Marquardt method (e.g., Moré (2006)), based on 12 that is specified in Alg. 1. Empirically, we observe Alg. 1 to converge after much fewer iterations than gradient descent algorithms (see § B.1 for more implementation details and experiments).
@@ 3.4 Extensions
where $c>0$ is a constant and $\tau>0$ is a temperature parameter. The smaller we choose $\tau$, the better the approximation in 13 becomes. The function $g$ corresponds to a conditional normalizing flow that was trained on the logits output by a classifier, obtained either by regressing on $\mathbf{x}_{\text{pa}(i)}$ or $x_{\text{img}}$ (such as the MNIST image in Fig. 1). We see that $f_{i}$ indeed fulfills the conditions of being both invertible and differentiable in $U_{i}$, thus enabling the application of DeepBC.
@@ 3.5 Properties in the Context of Counterfactual Explanations
    Causal Compliance. By construction, DeepBC generates counterfactuals that adhere to causal laws $(f_{1},f_{2},...,f_{n})$ by ensuring the preservation of these laws during the generation process: It delineates similarity between data points $\mathbf{x},\mathbf{x}^{\prime}$ in terms of their latent representations $\mathbf{u},\mathbf{u}^{\prime}$ that correspond to the pull-back through the causal laws as encoded by the reduced form, i.e., $\mathbf{u}=\mathbf{F}^{-1}(\mathbf{x})$ and $\mathbf{u}^{\prime}=\mathbf{F}^{-1}(\mathbf{x}^{\prime})$. This way, the laws are guaranteed to be retained in the generated counterfactual. This contrasts counterfactual explanations methods, which do not model the true causal mechanisms explicitly (see Fig. 6 for an illustrative example).
@@
    Versatility. DeepBC naturally supports dealing with complex causal relationships between multiple variables that are potentially high dimensional (e.g., images or scalar attributes). This goes beyond the instance-label setup 2 presented in § 2.4, and thus naturally supports flexible choices of antecedent variables (Fig. 12 (a)). This contrasts counterfactual explanation methods that are typically limited to antecedents with respect to the output of a predictive model. DeepBC further allows for sampling (§ 3.1.1) and varying the distance functions $d_{i}$ in 5 to obtain counterfactuals with different properties, such as noise preservation (Fig. 11 in § D.1) or sparsity (Fig. 6).
@@
    Modularity. The structural equations or mechanisms $(f_{1},f_{2},...,f_{n})$ encoding the causal relations between variables may change across different domains, for example, as the result of interventions or environmental changes. It has been postulated that such changes tend to manifest sparsely, meaning that only a few of the modules $f_{i}$ change at a time, while the others remain fixed (Schölkopf et al., 2021; Perry et al., 2022). By explicitly modeling the individual structural equations (as deep generative components), DeepBC exhibits a natural modularity. As a result, adapting to new domains only requires adjusting those components which undergo a domain shift while all other modules can be re-used. This contrasts with counterfactual explanation methods 2, which do not incorporate such replaceable modules and thus require relearning the entire model to handle domain shifts.

### v4

@@ 1 Introduction
The present work further serves as a bridge between causal modeling and practical methods in the field of counterfactual explanations (Wachter et al., 2017; Beckers, 2022). As a causally grounded approach applicable to high-dimensional data, our method fills a gap in the existing literature between non-causal explanation tools, built for complex data such as images (e.g., Goyal et al., 2019; Boreiko et al., 2022), and causal methods that have only been applied to simple (assuming additive noise), low-dimensional settings (Bynum et al., 2024; von Kügelgen et al., 2023b; Crupi et al., 2022).
@@ 2.1 Structural Causal Models
Let $\mathbf{X}=(X_{1},X_{2},...,X_{n})$ be a collection of potentially high-dimensional observable “endogenous” random variables. For instance, these variables could be high-dimensional objects such as images (e.g., the MNIST image in Fig. 1) or scalar feature variables (such as $t$ and $i$ in Fig. 1). The causal relationships among the $X_{i}$ are specified by a directed acyclic graph $G$ that is known. A structural causal model (Pearl, 2009) is characterized by a collection of structural equations $X_{i}\leftarrow f_{i}(\mathbf{X}_{\text{pa}(i)},U_{i})$, for $i=1,2,...,n,$ where $\mathbf{X}_{\text{pa}(i)}$ are the causal parents of $X_{i}$ as specified by $G$ and $\mathbf{U}=(U_{1},U_{2},...,U_{n})$ are exogenous latent variables [^1] We note that by definition, the exogenous variables are not causally related to each other.. The acyclicity of $G$ ensures that for all $i$, we can recursively solve for $X_{i}$ to obtain a deterministic expression in terms of $\mathbf{U}$. Thus, there exists a unique function that maps $\mathbf{U}$ to $\mathbf{X}$, which we denote by $\mathbf{F}$,
@@ 2.2 Deep Invertible Structural Causal Models
In this work, we make the simplifying assumption that $f_{i}(\mathbf{x}_{\text{pa}(i)},\,\cdot\,)$ is invertible for any fixed $\mathbf{x}_{\text{pa}(i)}$222also known as bijective generation mechanism (see, e.g., Nasr-Esfahany et al., 2023). such that we can write
@@
$\mathbf{U}=\mathbf{F}^{-1}(\mathbf{X}).$ (2)
@@
(Rezende & Mohamed, 2015; Winkler et al., 2019) are constructed as a composition of invertible functions, hence rendering the entire function $f_{i}$ invertible in $u_{i}$. In addition, they are chosen such that the determinant of the Jacobian can be computed efficiently. These two attributes facilitate efficient training of $f_{i}$ via maximum likelihood.
@@ 2.3 Interventional and Backtracking Counterfactuals
leave all structural assignments unchanged. In order to set the antecedent $\mathbf{x}^{*}_{S}\neq\mathbf{x}_{S}$ true, they trace differences to the factual realization back to (ideally small) changes in the latent variables $\mathbf{U}$. These modified latent variables are represented by a new collection of variables $\mathbf{U}^{*}$ that depend on $\mathbf{U}$ via a given backtracking conditional $p^{B}(\mathbf{u}^{\prime}\,|\,\mathbf{u})$ (red in Fig. 2(von Kügelgen et al., 2023b), which represents a probability density for computing similarity between $\mathbf{u}$ and $\mathbf{u}^{\prime}$ and which we assume to be decomposable, or factorized:
@@
$p^{B}_{i}(u^{\prime}_{i}\,|\,u_{i})\;\propto\;\text{exp}\left\{-d_{i}(u^{\prime}_{i},u_{i})\right\},$ (3)
where the $d_{i}$ are differentiable with respect to $u^{\prime}_{i}$. The functions $d_{i}(u^{\prime}_{i},u_{i})$ can be interpreted as distances that penalize deviations of the counterfactual latent variables to their factual realizations.
Backtracking counterfactuals fulfill an intuitive notion of causal compliance, which we will elaborate on the example shown in Fig. 2. To simplify the exposition, we only consider the modes (see § 3.1.2) of the backtracking variables (denoted by $u^{*}_{1}$, $u^{*}_{2}$, $u^{*}_{3}$, $x^{*}_{1}$, $x^{*}_{2}$ and $x^{*}_{3}$) and note that only the latent variables $U^{*}_{1}$ and $U^{*}_{2}$, which are upstream (in the causal graph) of the antecedent variable $X^{*}_{2}$, can contribute in realizing $x^{*}_{2}\neq x_{2}$. Thus, we generally have $u^{*}_{1}\neq u_{1}$ and $u^{*}_{2}\neq u_{2}$, i.e., the counterfactual modes differ from the factual realizations. The downstream latent variable $U^{*}_{3}$, in contrast, does not have a causal influence on $X^{*}_{2}$. In order to minimize the distance $d_{3}(u^{*}_{3},u_{3})$, $u^{*}_{3}$ is thus left unchanged from the factual $u_{3}$, i.e., $u^{*}_{3}=u_{3}$. The inequality $x^{*}_{3}\neq x_{3}$ is then solely a consequence of the downstream effect of the antecedent $x^{*}_{2}\neq x_{2}$. We will demonstrate these properties experimentally in Section § 4.
@@ 2.4 Counterfactual Explanations
where $d_{o}$ represents a distance function between observed variables. In the present work, we generally refer to methods implementing a variant of 4 as counterfactual explanations. We stress that $\hat{Y}$ is the prediction of a model ($f_{\hat{Y}}$) and thus always an effect of $X$. In general, the structural assignment $f_{\hat{Y}}$ is different from $f_{Y}$ (the assignment of the true variable $Y$ that is not predicted). For instance, $Y$ might be a cause of $X$ or might be confounded with $X$. We revisit the difference between $Y$ and $\hat{Y}$ in Section § 3.2.
@@ 3.1.1 Objective for Stochastic DeepBC
We sample from a distribution over counterfactuals $\mathbf{X}^{*}\,|\,\mathbf{x}^{*}_{S},\mathbf{x}$ for the factual realization $\mathbf{x}$, antecedent $\mathbf{x}^{*}_{S}$ and (known) backtracking conditional $p^{B}$. [^3] In contrast to the reduced form, the backtracking conditional is not learned from data, but must be specified explicitly. This distribution is characterized by the density
@@
where $\delta_{\mathbf{x}^{*}_{S}}(\,\cdot\,)$ refers to the dirac delta at $\mathbf{x}^{*}_{S}$. Intuitively, we can understand this distribution as describing counterfactuals that are likely given $\mathbf{x}$ in terms of latent components ($p^{B}$), while fullfilling the constraint of being compliant both with the antecedent $\mathbf{x}^{*}_{S}$ ($\delta_{\mathbf{x}^{*}_{S}}$) and with the causal laws $\mathbf{F}$. We further note that 5 is equivalent to a sampling problem within the structured latent space, i.e.,
@@
The distribution of $\mathbf{X}^{*}\,|\,\mathbf{x}^{*}_{S},\mathbf{x}$ in 5 is obtained as the push-forward of 6 by $\mathbf{F}$, for $\mathbf{x}=\mathbf{F}(\mathbf{u})$.
@@ 3.1.2 Objective for Mode DeepBC
We compute the mode of $p(\mathbf{x}^{\prime}\,|\,\mathbf{x}^{*}_{S},\mathbf{x})$ in 5, i.e., a single “most likely” counterfactual $\mathbf{x}^{*}$ for the factual realization $\mathbf{x}$ as a solution to the following constrained optimization problem:
@@
Intuitively, we can understand this optimization as finding a solution $\mathbf{x}^{*}$ that is close to the factual realization $\mathbf{x}$ in terms of its latent components. This situation is visualized on the Morpho-MNIST example in Fig. 1. We further note that 7 is equivalent to an optimization problem within the structured latent space,
@@
We obtain the solution of 7 by inserting the solution of 8 into $\mathbf{F}$.
@@ 3.2 Relation to Counterfactual Explanations
We can recover counterfactual explanations (§ 2.4) as a special form of mode DeepBC 7. To this end, we assume access to two variables with the following structural equations
@@
where we note that $\hat{Y}$ is not subject to additional randomness $U_{\hat{Y}}$. In this specific case, we observe that the mode DeepBC optimization problem 7 reduces to
@@
which can be interpreted as an instance of 4, where distance is measured in an unstructured latent space, governed by $f_{X}$. Under the assumption that $f_{X}$ is modeled as a deep invertible generative model (§ 2.2), we the refer to 10 as deep counterfactual explanation. For example, $x$ could be a high-dimensional image, $f_{X}$ an (unconditional) variational autoencoder and $y$ a label of the image. [^4] Typically, such methods do not explicitly model other variables besides $x$ and $y$. From this viewpoint, we can interpret DeepBC as a general form of counterfactual explanations in two ways: Firstly, it accommodates non-deterministic relations among variables, taking into account the influence of noise on all variables. In the aforementioned instance 9, this can be modeled by $\hat{Y}\leftarrow f_{\hat{Y}}(X,U_{\hat{Y}})$ (demonstrated in Fig. 12 (b) of § D.2). Secondly, DeepBC accounts for multiple variables with more general underlying causal relationships. For example, there could be a third variable $Z$ related to both $X$ and $Y$ in 10 that could be modeled as well, as depicted in Fig. 3 (a). Rather than treating $Z$ as a new dimension in the predictor input $f_{\hat{Y}}$ in 9 (resulting in $f_{\hat{Y}}(X,Z)$, see Fig. 3 (b)), DeepBC allows for explicitly modeling the true causal relations between $Z$, $X$ and $Y$ (see Fig. 3 (a)) via (deep) structural equations, as outlined in Section § 2.1. The benefit of DeepBC is that it intends to answer queries regarding the true underlying variables (see query in Fig. 3 (a)) rather than being confined to model predictions that generally do not model the causal mechanisms according to the true relationships (see query in Fig. 3 (b)). We elaborate further on the benefits of our approach in Section § 3.5.
@@
In the next section (§ 3.3), we propose algorithms to generate backtracking counterfactuals in practice, based on the formulations presented in 5 and 7.
@@ 3.3 Algorithms
We rely on a penalty formulation to approximate 6 and 8 in order to account for the dirac measure and the constraint, respectively. Specifically, we consider the following energy (loss) function with respect to $\mathbf{u}^{\prime}$:
@@ 3.3.1 Algorithm for Stochastic DeepBC
and hence we can use 12 to generate approximate samples from the desired distribution 6. In practice, we apply an Euler-Maruyama discretization (Sauer, 2013) of 12. It is given as
@@ 3.3.2 Algorithm for Mode DeepBC
We approximately compute the mode of $\mathbf{X}^{*}\,|\,\mathbf{x}^{*}_{S},\mathbf{x}$ by directly minimizing the energy $\mathcal{L}(\mathbf{u}^{\prime};\,\mathbf{u},\mathbf{x}^{*}_{S})$ 11 with respect to $\mathbf{u}^{\prime}$, where $\mathbf{u}=\mathbf{F}^{-1}(\mathbf{x})$. Rather than performing gradient descent on the original objective, we empirically observe that using a first-order Taylor approximation of $\mathbf{F}_{S}$ at $\bar{\mathbf{u}}$ is beneficial when minimizing the distance
@@
where $\bar{\mathbf{u}}$ is the linearization point and $\mathbf{J}_{S}(\bar{\mathbf{u}})\coloneqq\nabla_{\mathbf{u}}\mathbf{F}_{S}(\bar{\mathbf{u}})^{\top}$ denotes the Jacobian matrix. As a result of this approximation, 11 is a convex quadratic function in $\mathbf{u}^{\prime}$ and can therefore be solved for its minimum $\hat{\mathbf{u}}^{*}$ in closed form:
@@
and $\mathbf{W}\coloneqq\text{diag}(w_{i})$ is a diagonal matrix containing the distance weights $w_{i}$. By default, we set $w_{i}=1$, for all $i$. A different choice of weights may be useful to encode the notion that certain noise variables are more stable (where large $w_{i}$ corresponds to a high stability) due to application-specific requirements. A detailed derivation of 14 is provided in § A.3.
Solving 14 once, starting from the initial condition $\bar{\mathbf{u}}=\mathbf{u}$, does not accurately fulfill the constraint due to the constraint linearization, except for special cases. We thus apply an iterative algorithm similar to the Levenberg-Marquardt method (e.g., Moré (2006)), based on 14 that is specified in Alg. 1. Empirically, we observe Alg. 1 to converge after much fewer iterations than gradient descent algorithms (see § B.1 for more implementation details and experiments).
@@ 3.4 Extensions
where $c>0$ is a constant and $\tau>0$ is a temperature parameter. The smaller we choose $\tau$, the better the approximation in 15 becomes. The function $g$ corresponds to a conditional normalizing flow that was trained on the logits output by a classifier, obtained either by regressing on $\mathbf{x}_{\text{pa}(i)}$ or $x_{\text{img}}$ (such as the MNIST image in Fig. 1). We see that $f_{i}$ indeed fulfills the conditions of being both invertible and differentiable in $U_{i}$, thus enabling the application of DeepBC.
@@
Other distance functions.
In general, any combination of differentiable distance functions $d_{i}$ 3 can be used when applying gradient descent-based methods for mode DeepBC.
@@ 3.5 Properties in the Context of Counterfactual Explanations
    Causal Compliance. By construction, DeepBC generates counterfactuals that adhere to causal laws $(f_{1},f_{2},...,f_{n})$ by ensuring the preservation of these laws during the generation process: It delineates similarity between data points $\mathbf{x},\mathbf{x}^{\prime}$ in terms of their latent representations $\mathbf{u},\mathbf{u}^{\prime}$ that correspond to the pull-back through the causal laws as encoded by the reduced form, i.e., $\mathbf{u}=\mathbf{F}^{-1}(\mathbf{x})$ and $\mathbf{u}^{\prime}=\mathbf{F}^{-1}(\mathbf{x}^{\prime})$. This way, the laws are guaranteed to be retained in the generated counterfactuals, independent of the used backtracking conditional. This contrasts counterfactual explanations methods 4, which do not model the true causal mechanisms explicitly and thus may violate causal relationships (see Fig. 4 (b) and Fig. 6).
@@
    Versatility. DeepBC naturally supports dealing with complex causal relationships between multiple variables that are potentially high dimensional (e.g., images or scalar attributes). This goes beyond the instance-label setup 4 presented in § 2.4, and thus naturally supports flexible choices of antecedent variables (Fig. 12 (a)). This contrasts counterfactual explanation methods that are typically limited to antecedents with respect to one single output of a predictive model. DeepBC further allows for sampling (§ 3.1.1) and varying the distance functions $d_{i}$ in 7 to obtain counterfactuals with different properties, such as noise preservation (Fig. 11 in § D.1) or sparsity (Fig. 6).
@@
    Modularity. The structural equations or mechanisms $(f_{1},f_{2},...,f_{n})$ encoding the causal relations between variables may change across different domains, for example, as the result of interventions or environmental changes. It has been postulated that such changes tend to manifest sparsely, meaning that only a few of the modules $f_{i}$ change at a time, while the others remain fixed (Schölkopf et al., 2021; Perry et al., 2022). By explicitly modeling the individual structural equations (as deep generative components), DeepBC exhibits a natural modularity. As a result, adapting to new domains only requires adjusting those components which undergo a domain shift while all other modules can be re-used. This contrasts with counterfactual explanation methods 4, which do not incorporate such replaceable modules and thus require relearning the entire model to handle domain shifts.